OpenStudy (anonymous):

What is the value of x 4x + 8y = 8 x + 2y = 2 0 2 There is no x value as there is no solution to this system x can be any value as there are infinitely many solutions to this system

OpenStudy (skullpatrol):

"It is better to guide the user to an answer rather than doing all the work yourself. We discourage giving out answers without letting the asker at least take some part in the analysis - the aim of this site is to teach."

OpenStudy (anonymous):

I have tried answering this question 3 times.

OpenStudy (anonymous):

use elimination method

OpenStudy (anonymous):

Okay thank you. : )

OpenStudy (anonymous):

ur welcome

OpenStudy (raden):

if given equation linear 2 variables : a1x + b1y = c1 a2x + b2y = c2 just evaluate the proportion of : if a1/a2 = b1/b2 = c1/c2 then (x,y) can be any value as there are infinitely many solutions to this system

OpenStudy (anonymous):

Okay thx. : )

OpenStudy (raden):

glad to help ;)